### v1

# Massive Dirac-Pauli physics in lead-halide perovskites

Abstract
In standard quantum electrodynamics (QED), the so-called non-minimal (Pauli) coupling is suppressed for elementary particles and has no physical implications. Here, we show that the Pauli term naturally appears in a known family of Dirac materials – the lead-halide perovskites, suggesting a novel playground for the study of analogue QED effects. We outline measurable manifestations of the Pauli term in the phenomena pertaining to (i) the Klein paradox and (ii) relativistic corrections to bound states. In particular, we demonstrate that the binding energy of an electron in the vicinity of a positively charged defect is noticeably decreased due to the polarizability of lead ions and the appearance of a Darwin-like term. Our study adds to understanding of quantum phenomena in lead-halide perovskites, and paves the way for tabletop simulations of analogue Dirac-Pauli equations.
I Introduction
The band structure of certain condensed matter systems (Dirac materials) resembles the energy spectrum of the single-body relativistic problem [1]. Arguably, the best known system of this type is graphene whose massless Dirac-like band structure leads to a number of remarkable physical phenomena that were previously discussed mainly in the context of relativistic high-energy physics [2]. Topological insulators and Weyl semimetals are among the other celebrated examples [3, 4, 5]. These materials have become the tabletop experimental platforms for visualizing the extreme relativistic phenomena that obey conventional particle theories, such as QED. The natural question that may arise in this context is whether the analog Dirac materials can help to study conceivable high-energy effects that fall beyond the Standard Model of elementary particles.
The coupling of Dirac materials to electromagnetic fields is typically given via a minimal substitution: $p^{\nu}\rightarrow p^{\nu}-qA^{\nu}$, where $p^{\nu}$ and $A^{\nu}=(\phi,\bm{A})$ are the usual four-momentum and electromagnetic four-potential, respectively [6]. However, Lorentz and gauge invariance allow in general for additional ‘non-minimal’ couplings to electromagnetic fields (see, e.g., Ref. [7]). The first such extension of the Dirac equation was suggested by Pauli [8] who introduced what is now known as the Dirac-Pauli equation [9]
$\left(\gamma^{\nu}(p_{\nu}-qA_{\nu})-\frac{i\mu^{\prime}}{4}[\gamma_{\mu},\gamma_{\nu}]F^{\mu\nu}-m\right)\psi=0.$ (1)
Here, $\gamma^{\nu}$ are the Dirac matrices, $F^{\mu\nu}$ is the electromagnetic field tensor, $m$ is the rest mass of the Dirac particle and $\mu^{\prime}$ is its anomalous moment. The experimental data on magnetic moments of elementary particles suggest that $\mu^{\prime}$ is negligibly small [10]. It can even be assumed that $\mu^{\prime}=0$ at the current state of our understanding of quantum electrodynamics  [^1] Recently, it was proposed that the Pauli term can be important in the context of the Landau pole [38, 39]. This however does not change the fact that the value of $\mu^{\prime}$ must be negligibly small.. Therefore, the Pauli term is typically neglected with rare exceptions of effective quantum field theories that aim to include polarizabilities of composite particles such as neutrons and protons [12, 13, 14] in a phenomenological way.
Here, we show that the Pauli term naturally appears in the low-energy description of lead-halide perovskites (LHPs), making it (to the best of our knowledge) the first known Dirac-Pauli semiconductor. Lead-halide perovskites is a family of advanced energy materials [15] with exceptional optoelectronic properties [16, 17]. Charge carriers in LHPs can be described by the massive Dirac equation [18, 19] with a peculiarity that their coupling to electromagnetic fields requires an explicit inclusion of the high polarizability of Pb2+ ions [20, 21]. Here, we first illustrate that this non-standard matter-field interaction is actually the Pauli coupling with the (effective) atomic dipole polarizability in place of the anomalous moment, $\mu^{\prime}$. Further, to facilitate future theoretical and experimental studies, we outline a few effects that the Dirac-Pauli physics can have on the properties of LHPs. In particular, we consider scattering off a step potential and binding of charge carriers to local defects, such as vacancies and interstitials.
II Results
[…]
$\begin{split}|\!\!\Uparrow\uparrow\rangle=-\left(|p_{z}\rangle|\!\!\uparrow\rangle+\left(|p_{x}\rangle+i|p_{y}\rangle\right)|\!\!\downarrow\rangle\right)/\sqrt{3},\\ |\!\!\Uparrow\downarrow\rangle=\left(|p_{z}\rangle|\!\!\downarrow\rangle-\left(|p_{x}\rangle-i|p_{y}\rangle\right)|\!\!\downarrow\rangle\right)/\sqrt{3},\\ |\!\!\Downarrow\uparrow\rangle=|s\rangle|\!\!\uparrow\rangle,\quad\quad\quad|\!\!\Downarrow\downarrow\rangle=|s\rangle|\!\!\downarrow\rangle.\end{split}$ (2)
[…]
In the absence of external fields, the low-energy spectrum of LHPs is determined by diagonalizing the Hamiltonian ($\hbar=c=1$[18]:
$H=\Delta(\bm{k})\tau_{3}\otimes\mathbb{I}+2at\ \tau_{2}\otimes\bm{\sigma}\cdot\bm{k},$ (3)
where $\bm{\sigma}$ is a vector of Pauli matrices, so that $\bm{\sigma}\cdot\bm{k}=\sum_{i}\sigma_{i}k_{i}$. $\Delta(\bm{k})=\frac{1}{2}\bigg{(}\Delta+\frac{t_{3}a^{2}|\bm{k}|^{2}}{2}\bigg{)}$, $\bm{k}$ is the quasi-momentum counted from the high-symmetry $R$-point relevant for low-energy physics of LHPs, $\Delta$ is the bandgap at the $R$-point, $a$ is the lattice constant, $t_{3}$ and $t$ are further $\bm{k}\cdot\bm{p}$ parameters of the Hamiltonian. They can be estimated by fitting to numerical calculations or experimental data. For example, for methylammonium lead bromide (MAPbBr3) these parameters are found to be: $t\simeq 0.6$eV, $t_{3}\simeq 0.9$eV, $a\simeq 0.586$nm, $\Delta\simeq 2.3$eV [19, 21].
It was recently shown both experimentally and theoretically that in order to account for the full effect of the external electric field $\bm{E}$ on LHPs, one should employ not only the minimal coupling $\bm{\tilde{k}}=\bm{k}-q\bm{A}$ (where $q$ is the charge of an electron) in Eq. (3) but also an additional term $\mu\ \tau_{1}\otimes\bm{\sigma}\cdot\bm{E}$, dubbed ‘spin-electric’ in Refs. [20, 21] because it couples the quasi-spin degree of freedom to an external electric field. The non-minimal coupling term accounts for the local effects of the electric field on the lattice, in particular, on the highly polarizable Pb2+. Therefore, we interpret $\mu$ as coming from the effective atomic polarizability of lead ions. For MAPbBr3, the experiment demonstrates that the magnitude of this effect corresponds to $\mu\simeq-0.3|q|a$ [21].
A close examination reveals that the operator $\mu\ \tau_{1}\otimes\bm{\sigma}\cdot\bm{E}$ is nothing else but the electric part of the Pauli term in Eq. (1). Indeed, the Hamiltonian in Eq. (3) amended with the spin-electric term can be written as:
$H=\Delta(\bm{\tilde{k}})\beta+2at\bm{\alpha}\cdot\bm{\tilde{k}}+i\mu\bm{\gamma}\cdot\bm{E},$ (4)
where $\beta=\tau_{3}\otimes\mathbb{I}$, $\bm{\alpha}=\tau_{2}\otimes\bm{\sigma}$, and $\bm{\gamma}=\beta\bm{\alpha}$. As the matrices $\gamma^{\nu}=\{\beta,\bm{\gamma}\}$ satisfy the Dirac algebra $\{\gamma^{\mu},\gamma^{\rho}\}=\mathrm{diag}(2,-2,-2,-2)$, the first two terms in Eq. (4) lead to the Dirac equation in Eq. (1) with $\mu^{\prime}=0$ and $t_{3}=0$. [The ‘standard (Dirac) representation’ of the gamma matrices [9] is obtained by the unitary transformation $U^{\dagger}\bm{\alpha}U$ where $U=e^{-i\tau_{3}\pi/4}$, which corresponds to a $\pi/2$-rotation in the $\tau$-subspace.] Note that $t_{3}\neq 0$ is beyond the Dirac equation, as is common for gapped Dirac materials [25]. We note that while this term does not break the algebraic structure of the Dirac equation its presence leads to measurable consequences, e.g., in the determination of the energy of a bound state in an impurity potential (see below).
Finally, in the absence of magnetic fields $[\gamma^{\mu},\gamma^{\nu}]F^{\mu\nu}=4\bm{\gamma}\cdot{\bm{E}}$, establishing equivalence between Eqs. (1) and (4) with $\mu^{\prime}=\mu$. Note that the magnetic part of the Pauli term $[\gamma^{\mu},\gamma^{\nu}]F^{\mu\nu}$ leads to an anomalous $\tau_{3}\otimes\bm{\sigma}\cdot\bm{B}$ coupling, which also enters the low-energy description of LHPs placed in the magnetic field $\bm{B}$ [20, 21]. However, the precise magnitude of this magnetic term in MAPbBr3 is currently undetermined, and therefore, we leave the corresponding investigation to future studies.
In what follows, we illustrate the physical implications of the Pauli term for LHPs. To this end, we solve the stationary Dirac-Pauli equation
$H\Psi=\mathcal{E}\Psi,$ (5)
where $\Psi$ is the 4-component Dirac spinor, $\Psi^{T}=(\psi_{\mathcal{\Uparrow}\uparrow},\psi_{\mathcal{\Uparrow}\downarrow},\psi_{\mathcal{\Downarrow}\uparrow},\psi_{\mathcal{\Downarrow}\downarrow})$, and $\mathcal{E}$ is the energy. [Here, the first (last) two elements of the spinor are typically referred to as electron (hole) components.] Specifically, we solve Eq. (5) with the Dirac-Pauli Hamiltonian for two standard physical problems: scattering of free charge carriers off a potential step (Klein paradox; see Fig. 1(a)) and determination of the energy of bound states in the vicinity of an attractive potential.
Klein problem in LHPs. Klein paradox refers to a counter-intuitive behavior of scattering amplitudes off a step potential whose height is larger than twice the particle’s rest mass. In the first quantization picture, the paradox manifests itself in the seemingly unphysical values for the reflection and transmission coefficients. In particular, the Klein’s solution suggests that the transmitted current does not vanish even when the height of the barrier approaches infinity [26, 27] (for more details see Refs. [28, 29, 30]).
Figure 1: (a) Schematic of the Klein problem: Electrons of energy $\mathcal{E}$ and momentum $k$ are incident on a potential barrier of height $V_{0}$. If the value of $V_{0}$ is larger than the energy gap, $\Delta$, between the conduction and valence bands, then Klein paradox occurs – particles with momentum $K$ appear in the high potential region. (b) Reflection coefficient as a function of the (dimensionless) barrier height with and without Pauli coupling assuming that $\mathcal{E}=0.01\,\text{eV}$. The other parameters correspond to MAPbBr3 with $\mu<0$. For $\mu>0$, only the sign of $\mu$ is changed in comparison to the case with $\mu<0$.
Direct experimental observation of this physics in quantum electrodynamics would require access to extremely high electric fields not available in modern laboratories  [^2] The exception is electron beams in aligned crystals, which however can only be used for indirect studies of Klein paradox [40]. This motivates searches for tabletop set-ups that would realize Dirac’s equation, such as gapless Dirac materials [32, 33] or photonic superlattices [34]. The most direct analogy to the Klein paradox in its original formulation however requires working with gapped Dirac semiconductors such as LHPs. The corresponding bandgap $\Delta$ then would play the role of a pair-creation energy $2mc^{2}$ in QED, where $m$ is the mass of the particle that is being created. As the gap $\Delta$ can be about six orders of magnitude lower than $2mc^{2}$ the corresponding effects in Dirac semiconductors are much more accessible as compared to QED. Nevertheless, it should be noted that the very existence of an energy gap of a few eV may already result in a considerable experimental obstacle for the observation of Klein paradox in LHPs. Still, we find it fitting to start the discussion of the effect of the Pauli term by considering the corresponding solution of the Dirac equation.
For simplicity, we assume that $t_{3}=0$ here. Similar to the original formulation of the problem by Klein, we consider the one-dimensional potential
$\phi(x)=V_{0}\Theta(x),$ (6)
where $\Theta(x)$ is the Heaviside step function, and $V_{0}$ is the strength of the potential. For $\mu\neq 0$, the associated electric field $E_{x}=-\partial\phi/\partial x=-V_{0}\delta(x)$ enters the Dirac equation via the Pauli term. This leads to a term with a $\delta$-function in Eq. (4), giving rise to a discontinuity in the spinor components at $x=0$, which can be treated as a specific boundary condition.
We solve the Dirac-Pauli equation for each spinor component and arrive at the following expressions for the reflection and transmission coefficients (see Suppl. Note 1):
$R=\bigg{|}\frac{r-1}{r+1}\bigg{|}^{2},\qquad T=\frac{\mathrm{Im}\Lambda}{\lambda}\frac{4|r|^{2}}{\big{|}r+1\big{|}^{2}}e^{\frac{\mu V_{0}}{at}},$ (7)
where $\lambda=4iatk/(\Delta+2\mathcal{E})$ and $\Lambda=4iatK/(\Delta+2\mathcal{E}-2qV_{0})$, $r=\lambda e^{-\frac{\mu V_{0}}{at}}/\Lambda$ and $k$ and $K$ are the momenta of the charge carrier for $x<0$ and $x>0$. For $\mu\to 0$, these expressions turn into the original Klein solution [30]. Non-vanishing values of $\mu$ lead an exponential dependence on $V_{0}$ in Eq. (7). Depending on the sign of $\mu$, the exponent $e^{-\frac{\mu V_{0}}{at}}$ that appears in the parameter $r$ is either increasing or decreasing, leaving marked imprints on the transmission and reflection coefficients.
When $\mu<0$ as in LHPs, the Klein paradox is modified in comparison to the case when there is no Pauli coupling. Curiously, the introduction of the Pauli term suppresses the anomalous ($R>1$) Klein reflectivity in the asymptotic limit of large barrier heights, $qV_{0}\rightarrow\infty$. Figure 1(b) illustrates this for the parameters of MAPbBr3. The behavior of the reflection coefficient for $\mu>0$ is even more intriguing as it features a resonance-like divergent behavior, which can be interpreted as a resonant enhancement of particle-antiparicle creation within the potential barrier. This might be indicative of a novel vacuum breakdown channel that has previously been overlooked (see Suppl. Note 2 for additional details). We conclude that for the either sign of $\mu$, Pauli coupling strongly affects Klein’s physics.
Bound states near static impurities in LHPs. In this subsection, we investigate how the bound states of an electron in the vicinity of a spherically symmetric local impurity potential, $\phi$, are modified by the Pauli term. To this end, we focus on the effective behavior of electrons by integrating out the hole degrees of freedom following the procedure of Ref. [35] (see Suppl. Note 3), which resembles Foldy-Wouthuysen transformation in QED [36].
By doing so, we frame the problem in the form of the Schrödinger equation for electrons
$\bigg{(}-\frac{\nabla^{2}}{2m_{e}^{*}}+q\phi(r)-\frac{q^{2}\phi(r)^{2}}{\Delta}+H_{P}\bigg{)}\psi(r)=\epsilon\psi(r),$ (8)
where $\psi^{T}=(\psi_{\mathcal{\Uparrow}\uparrow},\psi_{\mathcal{\Uparrow}\downarrow})$ is the electronic wavefunction; $\epsilon=\mathcal{E}-\Delta/2$ is the corresponding energy; $m_{e}^{*}=(8a^{2}t^{2}/\Delta+t_{3}a^{2}/2)^{-1}$ is the effective mass, note that the parameter $t_{3}$ enters only in the effective mass, otherwise its presence does not affect our results; $q\phi(r)-\frac{q^{2}\phi(r)^{2}}{\Delta}$ is the potential energy of an electron in the electric field generated by the impurity; $H_{P}$ is the sum of terms arising from Pauli coupling:
$H_{P}=\frac{\big{|}\mu\bm{E}(\bm{r})\big{|}^{2}}{\Delta}-\frac{2at\mu}{\Delta}{\bm{\nabla}}\cdot\bm{E}(\bm{r})-\frac{8at\mu}{\Delta}\frac{|\bm{E}(\bm{r})|}{r}\bm{S}\cdot\bm{L}.$ (9)
The first term in $H_{P}$ mimics the effect of polarizability of quasi-electrons. This term is non-vanishing even if $t=0$, implying that this polarizability of electrons inside lead-halide perovskites can be traced back to the atomic polarizability of the underlying lattice ions (primarily Pb2+), see also the Discussion section below. The $\bm{\nabla}\cdot\bm{E}(\bm{r})$ term is referred to as the Darwin term; in QED, this term is interpreted as coming from the jitter motion ($Zitterbewegung$) of a relativistic electron on the background of a non-uniform electric field [36]. The last term in $H_{P}$ describes the spin-orbit coupling (SOC).
If $\mu\neq 0$, SOC appears as $\sim\mu|\bm{E}(\bm{r})|\bm{S}\cdot\bm{L}/(r\sqrt{m_{e}^{*}\Delta})$ in the effective Schrödinger equation (8) of the Dirac-Pauli system. Hence, its influence on the states and energies is different when compared to a Dirac system of an electron in a given electric field. For example, in a hydrogen atom SOC has the paradigmatic form $\sim q|\bm{E}(\bm{r})|\bm{S}\cdot\bm{L}/(rm_{e}\Delta)$ [36]. We see that the above presented terms have identical functional dependence. However, their strengths are different with the ratio $\mu\sqrt{m_{e}^{*}\Delta}$. A similar conclusion can be reached also for the Darwin-like term.
Figure 2: (a) Ground state wavefunction with and without the Pauli term for $\frac{\mu}{aq}=0.3$. (b) Relative change in the binding energy of the electron due to the Pauli coupling plotted as a function of the binding energy with $H_{P}=0$. Note that the Pauli coupling leads to a decrease in the binding energy. Different curves are for different values of the effective dipole moment, $\mu$; the curve in the middle is calculated with the parameters of MAPbBr3.
To illustrate the effect of $H_{P}$ on the properties of bound states, we investigate the ground state of an electron interacting with a model Gaussian defect whose charge density is chosen as
$\rho(r)=\frac{Q}{(a\sqrt{\pi})^{3}}e^{-\frac{r^{2}}{a^{2}}},$ (10)
where $Q$ is the charge of the defect. The width of the Gaussian is determined by the lattice constant, $a$, to model a single-site defect of a semiconductor. We consider a defect with $Q>0$, which binds electrons. However, the hole-electron symmetry allows us to also easily understand what happens for $Q<0$ from our calculations.
The charge density produces a radial electric field given by
$E(r)=\frac{Q}{4\varepsilon\pi^{3/2}ar^{2}}\bigg{(}a\sqrt{\pi}\;\mathrm{erf}(r/a)-2re^{-\frac{r^{2}}{a^{2}}}\bigg{)},$ (11)
where $\varepsilon$ is the (dynamic) dielectric constant of the medium and $\mathrm{erf}$ is the standard error function. Once the electric field is known, we solve the Schrödinger equation (30) numerically by diagonalizing the corresponding matrix (see Suppl. Note 3). Specifically, we calculate the ground state energy of the system with and without $H_{P}$, and study the difference introduced by $H_{P}$.
For the ground state $L=0$, hence, the spin-orbit coupling term does not play a role in Eq. (31) and can be ignored. In other words, effects of SOC can only be seen in excited states. A study of the ground state allows us to demonstrate the effect of the first two terms in $H_{P}$: the polarizability and the Darwin terms. Note that in LHPs, the parameter $\mu$ is negative, which implies that the matrix $H_{P}$ is positive-definite, and hence, that the ground state energy is increased by the presence of the Pauli coupling.
We show our findings in Fig. 2. First, we plot the electronic density in Fig. 2(a). We see that the wave function becomes spatially broader as the Pauli term repels the electron. This change in the electronic density affects other observables as well, for example the binding energy $\epsilon_{B}>0$ of the electron-defect system. Indeed, Fig. 2(b) confirms that the Pauli term leads to a noticeable reduction of $\epsilon_{B}$. In the figure, we plot the relative decrease in binding energy:
$\delta=\frac{\epsilon_{B}(\mu=0)-\epsilon_{B}(\mu)}{\epsilon_{B}(\mu=0)}.$ (12)
We conclude that the Pauli coupling in lead-halide perovskite can have a measurable effect on the energies of electrons in the vicinity of defects. This effect scales with the parameter $\mu$, which depends on the chemical structure of the material. As we expect that $\mu$ can be affected by one’s choice of the halogen atom, further studies are needed to determine the lead-halide perovskite with the most pronounced effect of Pauli coupling for tabletop simulations of the corresponding physics.
III Discussion
It was noted by Foldy that the presence of the Pauli term implies finite polarizability of the particle in question [13]. Indeed, for a Hamiltonian of the form $H=\beta m+\bm{\alpha}\cdot\bm{k}+i\mu\bm{\gamma}\cdot\bm{E}$, the energy of the zero-momentum mode reads
$\epsilon^{k=0}=\sqrt{m^{2}+\mu^{2}E^{2}}\sim m+\frac{\mu^{2}E^{2}}{2m}.$ (13)
The energy shift quadratic in the electric field forces an interpretation in terms of the polarizability of the particle, which is not associated with any actual structural deformation. We interpret the Pauli term in LHPs in a similar manner: it describes an effective polarizability of an electron (and hole) quasiparticle in LHPs, which on the microscopic level can be traced back to the local atomic polarizability of the perovskite lattice, which is large due the presence of Pb2+ ions. In other words, the spin-electric term $\mu\ \tau_{1}\otimes\bm{\sigma}\cdot\bm{E}$ originates from a coupling between valence and conduction bands that is mediated by the electric field even in the limit $k=0$ due to the lattice polarizability. The fact that the value of $\mu$ was found to be considerable in LHPs [21] implies that quasi-electrons dressed by the high polarizability of lead ions obey the Dirac-Pauli equation.
To the best of our knowledge, LHP is the only known family of materials where the Pauli term cannot be neglected, providing a convenient testground to study intriguing theoretical concepts from high-energy physics. In this paper, we demonstrate some basic physical manifestations of Pauli coupling in the problems related to scattering off step potential and the spectrum of electronic bound states in the vicinity of defects. However, there are other scenarios where the effects of the Dirac-Pauli physics could be interesting to explore. For example, recent observations of the dynamical Schwinger effect [37] opens an opportunity to simulate analogues of relativistic strong-field effects, and to investigate (both theoretically and experimentally) the Pauli coupling in the presence of strong laser fields.